# Zero-shot Action Localization via the Confidence of Large Vision-Language Models

### v1

@@ Abstract
Precise action localization in untrimmed video is vital for fields such as professional sports and minimally invasive surgery, where the delineation of particular motions in recordings can dramatically enhance analysis. But in many cases, large scale datasets with video-label pairs for localization are unavailable, limiting the opportunity to fine-tune video-understanding models. Recent developments in large vision-language models (LVLM) address this need with impressive zero-shot capabilities in a variety of video understanding tasks. However, the adaptation of image-based LVLMs, with their powerful visual question answering capabilities, to action localization in long-form video is still relatively unexplored. To this end, we introduce a true ZEro-shot Action Localization method (ZEAL). Specifically, we leverage the built-in action knowledge of a large language model (LLM) to inflate actions into highly-detailed descriptions of the archetypal start and end of the action. These descriptions serve as queries to LVLM for generating frame-level confidence scores which can be aggregated to produce localization outputs. The simplicity and flexibility of our method lends it amenable to more capable LVLMs as they are developed, and we demonstrate remarkable results in zero-shot action localization on a challenging benchmark, without any training. Source code will be made available.
@@ 1 Introduction
More recently, several lines of work have attempted to address this issue in the zero-shot setting [9, 22, 34, 26], where not all action categories that are present in testing data are seen during training. The most common approach is localization in zero-shot with tuning, where frozen CLIP [27] vision and text encoders are combined with learnable prompt vectors [9] and modules [22, 26] to identify action intervals. However, these works are constrained to CLIP pretraining, which provides high-level semantic alignment between whole images and text. Although CLIP exhibits strong performance in cross-modal alignment tasks, AL using CLIP features alone largely ignores developments in more fine-grained alignment for local details. Recent advances in large vision-language models (LVLM) have demonstrated remarkable potential for dense visual reasoning tasks like grounded visual question answering [40, 41], GUI navigation [6] and other cross-modality tasks [33]. These single-image-based models have strong visual localization capabilities that enable intricate associations between visual elements and textual semantics.
@@
Although limited to single images, LVLM’s ability to answer visual questions compliments the the observation that we as humans naturally follow a similar process when seeking the boundaries of particular actions in videos. When presented with a novel action and tasked to find it in long-form video, there is a tendency to define salient portions of this action via language: "what does the start of a baseball swing look like?" "a bat angled upward, eyes on the ball, knees slightly bent" or "how do you know a basketball dunk is complete?" "person on two feet, ball on ground, person facing away from hoop". Humans then look for these defining moments in video at the frame-level and localize actions. Thus, we posit that action localization can be broken down into two steps: one where actions are deconstructed into their components similar in spirit to [10, 19] (i.e. how the start of the action should look, how the end of the action should look, etc.), and the second where visual-temporal inputs are parsed largely guided via language.
@@
In this work, we aim to leverage LVLMs to reflect this decomposition and reasoning process. Leveraging the knowledge of everyday actions from a LLM, we decompose generic action classes into precise descriptors of the beginning and the end of these activities. We use these descriptors as queries for an LVLM to parse visual information per frame and generate confidence scores. Specifically, we prompt LVLM to answer affirmative or negative to questions at the frame-level and extract the model’s confidence in answering these visual questions. The resulting scores serve as a proxy for the likelihood of certain orientations and semantics expected at the start of an activity and likewise the end of the activity. From these scores we produce candidate localization outputs by matching start and end times which coincide with the most confident frame-level scores. This framework is agnostic to choice of LLM for action decomposition in generating descriptors or queries, and to choice of LVLM for interpreting visual inputs and producing confidence scores.
@@
Our paper contributes 1) A novel framework leveraging LLM for action decomposition and LVLM for parsing visual inputs to perform true ZEro-shot Action Localization (ZEAL), 2) a new method for creating a surrogate for the confidence of LVLM in answering localization queries, and 3) without any task-specific training, we achieve competitive or state-of-the-art zero-shot results compared to previous training-based methods on THUMOS14, a representative and challenging AL benchmark of every-day human activities and sports, demonstrating our framework’s flexibility in localizing actions on long-form, untrimmed video.
@@ 3 Method
3.1 Generating Vision-Language Queries for Localization
The first stage of our framework involves generating queries for the LVLM to appropriately identify key markers of the actions in question for localization. To this end, we employ an LLM (e.g. language-only GPT-4 [24]) to decompose action classes into two questions, one seeking what the archetypal start of the action should look like (), and the other what the archetypal end of the action should look like (). Importantly, we enforce that these questions are directly "yes/no" answerable and prompt the LVLM in subsequent stages to only answer affirmative or negative. Prior work [20] has successfully demonstrated that the decomposition of class categories into semantic descriptors via mining of an LLM enhances image classification. We employ a similar idea here by extracting the built-in action knowledge of LLM to inflate action classes into descriptors that delineate the boundaries of these actions. This allows the LVLM to parse its visual inputs and subsequently localize actions. Moreover, we prompt the LLM to give a short description of the action () that will be used to construct a CLIP-based actionness score (see Section 3.2). Here is an example of a produced start and end query, along with a short description, for the “CliffDiving" action:
0cm-1cm
@@ 3.1 Generating Vision-Language Queries for Localization
If all action classes are known a priori, these queries need only be generated once, incurring minimal costs if using proprietary APIs even with large . We include examples of prompts and outputs in Appendix A.1.
@@ 3.2 Coarse Action Class Filtering for Relevancy
To reduce computational overhead introduced by repeated forward passes through the LVLM, we implement an instance-level filtering of action classes that are relevant to the untrimmed video. This is done primarily to reduce the search space for LVLM question answering. The instance-level filtering is implemented as follows: given a closed set of action classes, we sample frames from the video and use CLIP [37] to embed these frames. We collect the top-k action classes (cosine similarity w.r.t the image features) most likely to occur in a video containing these frames. Please see Appendix A.2 for more performance details of this component.
@@ 3.3 CLIP-based Actionness
We also utilize the short description produced by the LLM for a given action and measure the cosine similarity () to frame-level CLIP [27] features. Again, we use a CLIP vision tower for extracting frame-level features and the corresponding text tower to embed the description (see Appendix A.3). The cosine similarities are used when generating action interval proposals (see Section 3.5) to avoid creating intervals containing background frames.
@@ 3.4 Fine-grained Localization with LVLM Outputs
At the first step of generation, we extract the logits of the "yes" token () and the "no" token () and compute two scores (one for and ) for frame as . These scores serve as confidence scores of the LVLM in affirming the question about the current state of the activity (as defined by or ). This also results in two disparate distributions, one which peaks at the beginning of action instances in the untrimmed video, and one that peaks at the ends of these instances (see Figure 2).
@@ 3.5 Creating Action Interval Proposals
Our framework (see Figure 2) can leverage LLM and LVLM for action decomposition and visual perception with the ultimate goal of producing action localization outputs. With minimal hyperparameters ( and ), our proposed method does not require any parameter optimization or training.
4 Empirical Results
@@ 4.1 Datasets and Evaluation Settings
We report the experimental results of previous methods and our proposed framework across two different evaluation settings: (Zero-shot 75% | 25%) where 75% of action classes with labels appear during training, while the remaining 25% are only seen at inference and (Zero-shot 50% | 50%) where 50% of action classes appearing during training, and the remaining 50% at inference.
Performance in this work is measured by the standard mean average precision (mAP) at different intersection-over-union (IoU) thresholds as reported in prior art. For producing splits, we randomly sample 10 subsets of the action classes and report the average over these splits to guarantee statistical significance. Note that our method doesn’t use any of the classes in the train splits, but we report our results on the same randomly sampled test split as the other baselines.
THUMOS14 Setting Method OOD TF NTTO mAP@0.3 0.4 0.5 0.6 0.7 Avg Fully supervised ActionFormer [39] ✗ ✗ ✓ 82.1 77.8 71.0 59.4 43.9 66.8 Zero-shot 75% Seen, 25% Unseen EffPrompt [9] ✗ ✗ ✓ 39.7 31.6 23.0 14.9 7.5 23.3 STALE [22] ✗ ✗ ✓ 40.5 32.3 23.5 15.3 7.6 23.8 MMPrompt [10] ✗ ✗ ✓ 46.3 39.0 29.5 18.3 8.7 28.4 ZEETAD [26] ✗ ✗ ✓ 61.4 53.9 44.7 34.5 20.5 43.2 EffPrompt† ✓ ✗ ✓ 7.1 5.9 4.5 3.4 2.2 4.6 STALE† ✓ ✗ ✓ 0.5 0.3 0.2 0.2 0.2 0.3 T3AL0 [13] ✓ ✓ ✓ 11.1 6.5 3.2 1.5 0.6 4.6 T3AL ✓ ✓ ✗ 19.2 12.7 7.4 4.4 2.2 9.2 ZEAL-ca ✓ ✓ ✓ 25.4 15.7 9.2 4.1 1.3 11.2 ZEAL-ca✛ ✓ ✓ ✓ 31.4 19.9 11.4 5.1 1.8 13.9 ZEAL-ov ✓ ✓ ✓ 25.2 16.2 10.4 5.4 1.7 11.8 ZEAL-ov✛ ✓ ✓ ✓ 33.9 22.0 14.0 7.3 2.3 15.9 Zero-shot 50% Seen, 50% Unseen EffPrompt [9] ✗ ✗ ✓ 37.2 29.6 21.6 14.0 7.2 21.9 STALE [22] ✗ ✗ ✓ 38.3 30.7 21.2 13.8 7.0 22.2 MMPrompt [10] ✗ ✗ ✓ 42.3 34.7 25.8 16.2 7.5 25.3 ZEETAD [26] ✗ ✗ ✓ 45.2 38.8 30.8 22.5 13.7 30.2 EffPrompt† ✓ ✗ ✓ 5.4 4.4 3.5 2.7 1.9 3.6 STALE† ✓ ✗ ✓ 1.3 0.7 0.6 0.6 0.4 0.7 T3AL0 [13] ✓ ✓ ✓ 11.4 6.8 3.5 1.7 0.6 4.8 T3AL ✓ ✓ ✗ 20.7 14.3 8.9 5.3 2.7 10.4 ZEAL-ca ✓ ✓ ✓ 24.0 14.6 8.5 3.8 1.3 10.5 ZEAL-ca✛ (Ours) ✓ ✓ ✓ 30.3 19.1 10.9 4.9 1.7 13.4 ZEAL-ov ✓ ✓ ✓ 23.7 14.6 9.2 4.8 1.5 10.7 ZEAL-ov✛ ✓ ✓ ✓ 33.0 21.3 13.2 6.8 2.3 15.3
@@
Implementation details. In our work, we use GPT-4 [24] (gpt-4-0125-preview) as the LLM to produce vision-language queries for localization, and we employ two representative LVLMs that represent the dominant training designs of modern LVLM: we use CogAgent [6] as a highly-capable LVLM for its ability to reason over small elements and fine-grained details in images (due to training on a mixture of high and low resolution data), and we also use LLaVA-OneVision [43, 12] representing a class of LVLMs built on the LLaVA [17] framework (fine-tuning a visual projector to map visual inputs into the input space of LLM). In all of our empirical experiments, we find that and works best for the hyperparameters. We conduct all experiments on NVIDIA L40S or NVIDIA RTX A6000 GPUs, processing videos in parallel if multiple GPUs are available. Please see Appendix A.4 for more details.
@@ 4.2 Zero-shot Action Localization
Our method outperforms T3AL’s best baseline up to a 28.3% improvement in average mAP on the 75% | 25% split, and up to a 2.9% improvement on the 50% | 50% split, using only pretrained LVLM to affirm visual questions and generate soft-scores, and without parameter updates of any kind. If video-level labels are available (if the set of action classes for a given video is constrained), our method enjoys strong performance gains. We also note that using LLaVA-OneVision as the LVLM in our setup performs marginally better than using CogAgent as LVLM.
@@ 4.3 Ablations
In addition, we conduct ablations to investigate the contributions of different components of our framework. All ablations reported here are performed on the THUMOS14 benchmark (50% | 50% split) and assuming action classes for a given video are known.
CLIP for actionness scores. First, we ablate the backbone used for deriving actionness scores for intervals (as in Section 3.3). We test EVA-02-CLIP-L/14+ [31] and SigLIP-SoViT-400M/14 [37] as strong CLIP models for image and text embedding. We show downstream localization performance in Table 2 when using either backbone for actionness, and we find that SigLIP is the best performing CLIP backbone for our framework.
Effect of actionness scores for suppression. We also ablate the use of actionness scores via the tradeoff parameter. corresponds to where only is used to score intervals, while indicates that is ignored entirely (i.e. only boundary confidence scores contribute to an interval’s score calculation). Figure 4 depicts the effect on average mAP for varying values of the parameter, keeping all other components of the framework fixed. We observe that is the optimal value for the THUMOS14 benchmark, but also that average mAP is greater when than when . This implicates two important findings: 1) neither or alone is sufficient for assessing proposal likelihood (having highly confident boundary predictions and highly confident semantic conent within the interval is not mutually exclusive) and 2) actionness is more critical to performance than boundary confidence, but only by a small margin.
Limitations and Future Work. The primary limitation of the proposed framework is the computational overhead of inference with LVLM. This may impact applications where near real-time inference speeds are necessary for utility. Another limitation of this framework lies in the need to filter relevant action classes to a given sample. When the number of potential action classes is large, this can limit the performance of our framework relying on action-specific queries to LVLM. One could instead use LVLM to generate captions or action classes from key frames of the video itself to limit the size of the query space, and we leave this to future work.
In addition, the reliance of our method on the logits of LVLM requires public access to the model. Also, extracting the logits of the "yes" and "no" tokens as described in this work may not be the only way to assess confidence of LVLM in answering visual questions. Instead, one could inflate action classes with key phrases that describe different phases of the action and observe the enrichment of the phrase’s token logits at each frame. This strategy and other prompting strategies for LLM could result in more fine-grained localization outputs, but we leave this as a promising direction for future work.
5 Conclusion

### v2

@@ Abstract
Precise action localization in untrimmed video is vital for fields such as professional sports and minimally invasive surgery, where the delineation of particular motions in recordings can dramatically enhance analysis. But in many cases, large scale datasets with video-label pairs for localization are unavailable, limiting the opportunity to fine-tune video-understanding models. Recent developments in large vision-language models (LVLM) address this need with impressive zero-shot capabilities in a variety of video understanding tasks. However, the adaptation of LVLMs, with their powerful visual question answering capabilities, to zero-shot localization in long-form video is still relatively unexplored. To this end, we introduce a true ZEro-shot Action Localization method (ZEAL). Specifically, we leverage the built-in action knowledge of a large language model (LLM) to inflate actions into detailed descriptions of the archetypal start and end of the action. These descriptions serve as queries to LVLM for generating frame-level confidence scores which can be aggregated to produce localization outputs. The simplicity and flexibility of our method lends it amenable to more capable LVLMs as they are developed, and we demonstrate remarkable results in zero-shot action localization on a challenging benchmark, without any training. Our code is publicly available at https://github.com/josaklil-ai/zeal.
@@ 1 Introduction
More recently, several lines of work have attempted to address this issue in the zero-shot setting [8, 20, 32, 23], where not all action categories that are present in testing data are seen during training. The most common approach is localization in zero-shot with tuning, where frozen CLIP [24] vision and text encoders are combined with learnable prompt vectors [8] and modules [20, 23] to identify action intervals. However, these works are constrained to CLIP pretraining, which provides high-level semantic alignment between whole images and text. Although CLIP exhibits strong performance in cross-modal alignment tasks, AL using CLIP features alone largely ignores developments in more fine-grained alignment for local details. Recent advances in large vision-language models (LVLM) have demonstrated remarkable potential for dense visual reasoning tasks like grounded visual question answering [39, 38], GUI navigation [5] and other cross-modality tasks [31]. These multi-modal models have strong visual localization capabilities that enable intricate associations between visual elements and textual semantics.
@@
Although sometimes limited in multi-image or video comprehension, LVLM’s ability to answer basic visual questions compliments the the observation that we as humans naturally follow a similar process when seeking the boundaries of particular actions in videos. When presented with a novel action and tasked to find it in long-form video, there is a tendency to define salient portions of this action via language: “what does the start of a baseball swing look like?” “a bat angled upward, eyes on the ball, knees slightly bent” or “how do you know a basketball dunk is complete?” “person on two feet, ball on ground, person facing away from hoop”. Humans then look for these defining moments in video at the frame-level and localize actions. Thus, we posit that action localization can be broken down into two steps: one where actions are deconstructed into their components similar in spirit to [9, 18] (i.e. how the start of the action should look, how the end of the action should look, etc.), and the second where visual-temporal inputs are parsed largely guided via language.
@@
In this work, we aim to leverage LVLMs to reflect this decomposition and reasoning process. Leveraging the knowledge of everyday actions from a LLM, we decompose generic action classes into precise descriptors of the beginning and the end of these activities. We use these descriptors as queries for an LVLM to parse visual information per frame and generate confidence scores. Specifically, we prompt LVLM to answer affirmative or negative to questions at the frame-level and extract the model’s confidence in answering these visual questions. The resulting scores serve as a proxy for the likelihood of certain orientations and semantics expected at the start of an activity and likewise the end of the activity. From these scores we produce candidate localization outputs by matching start and end times which coincide with the most confident frame-level scores. This framework is agnostic to choice of LLM for action decomposition in generating descriptors or queries, and to choice of LVLM for interpreting visual inputs and producing confidence scores (see Fig. 1).
@@
Our paper contributes 1) A novel framework leveraging LLM for action decomposition and LVLM for parsing visual inputs to perform true ZEro-shot Action Localization (ZEAL), 2) a new method for creating a surrogate for the confidence of LVLM in answering localization queries, and 3) without any task-specific training, we achieve competitive or state-of-the-art zero-shot results compared to previous training-based methods on THUMOS14, a representative and challenging AL benchmark of every-day human activities and sports, demonstrating our framework’s flexibility in localizing actions in long-form, untrimmed video.
@@ 3 Method
3.1 Generating Vision-Language Queries for Action Localization
The first stage of our framework involves generating queries for the LVLM to appropriately identify key markers of the actions in question for localization. To this end, we employ an LLM (e.g. GPT-4o [30]) to decompose action classes into two questions, one seeking what the archetypal start of the action should look like (), and the other what the archetypal end of the action should look like (). Importantly, we enforce that these questions are directly “yes/no” answerable and prompt the LVLM in subsequent stages to only answer affirmative or negative. Prior work [19] has successfully demonstrated that the decomposition of class categories into semantic descriptors via mining of an LLM enhances image classification. We employ a similar idea here by extracting the built-in action knowledge of LLM to inflate action classes into descriptors that delineate the boundaries of these actions. This allows the LVLM to parse its visual inputs and subsequently localize actions. Moreover, we prompt the LLM to give a short description of the action () that will be used to construct a CLIP-based actionness score (see Section 3.2). Here is an example of a produced start and end query, along with a short description, for the “CliffDiving” action:
@@ 3.1 Generating Vision-Language Queries for Action Localization
If all action classes are known a priori, these queries need only be generated once, incurring minimal costs if using proprietary APIs even with large .
@@ 3.2 Coarse Action Class Filtering for Relevancy
To reduce computational overhead introduced by repeated forward passes through the LVLM, we implement an instance-level filtering of action classes that are relevant to the untrimmed video. This is done primarily to reduce the search space for LVLM question answering. The instance-level filtering is implemented as follows: given a closed set of action classes, we sample frames from the video and use CLIP [35] to embed these frames. We collect the top-k action classes (cosine similarity w.r.t image features) most likely to occur in a video containing these frames (see Section 4.2 for more details).
@@ 3.3 CLIP-based Actionness
We also utilize the short description produced by the LLM for a given action and measure the cosine similarity () to frame-level CLIP [24] features. Again, we use a CLIP vision tower for extracting frame-level features and the corresponding text tower to embed the description. The cosine similarities are used when generating action interval proposals (see Section 3.5) to avoid creating intervals containing background frames.
@@ 3.4 Fine-grained Localization with LVLM Outputs
At the first step of generation, we extract the logits of the ”yes” token () and the ”no” token () and compute two scores (one for and ) for frame as . These scores serve as confidence scores of the LVLM in affirming the question about the current state of the activity (as defined by or ). This also results in two disparate distributions, one which peaks at the beginning of action instances in the untrimmed video, and one that peaks at the ends of these instances (see Fig. 2 and Fig. 4).
@@ 3.5 Creating Action Interval Proposals
Our framework (see Fig. 2) can leverage LLM and LVLM for action decomposition and visual perception with the ultimate goal of producing action localization outputs. With minimal hyperparameters ( and ), our proposed method does not require any parameter optimization or training.
4 Results
@@ 4.1 Datasets and Evaluation Settings
We report the experimental results of previous methods and our proposed framework across two different evaluation settings: (Zero-shot 75% — 25%) where 75% of action classes with labels appear during training, while the remaining 25% are only seen at inference and (Zero-shot 50% — 50%) where 50% of action classes appearing during training, and the remaining 50% at inference.
Performance in this work is measured by the standard mean average precision (mAP) at different intersection-over-union (IoU) thresholds as reported in prior art. For producing splits, we randomly sample 10 subsets of the action classes and report the average over these splits to guarantee statistical significance. Note that our method doesn’t use any of the classes in the train splits.
Implementation details. In our work, we use GPT-4o [30] (gpt-4o-2024-11-20) as the LLM to produce vision-language queries for localization, and we employ two representative LVLMs that represent the dominant training designs of modern LVLM: we use Pixtral (12B) [1] as a highly-capable LVLM for its ability to reason over fine-grained details in images (due to a new vision encoder trained from scratch to natively handle variable image sizes and aspect ratios), and we also use LLaVA-OneVision (7B) [41, 11] representing a class of LVLMs built on the LLaVA [16] framework (fine-tuning a visual projector to map visual inputs into the input space of LLM). In our empirical experiments, we find that and works best for the hyperparameters. We conduct all experiments on NVIDIA L40S or NVIDIA RTX A6000 GPUs, processing videos in parallel if multiple GPUs are available.
THUMOS14 | ||||||||||
Setting | Method | OOD | TF | NTTO | mAP@0.3 | 0.4 | 0.5 | 0.6 | 0.7 | Avg |
Fully supervised | ActionFormer [37] | ✗ | ✗ | ✓ | 82.1 | 77.8 | 71.0 | 59.4 | 43.9 | 66.8 |
Zero-shot 75% Seen, 25% Unseen | EffPrompt [8] | ✗ | ✗ | ✓ | 39.7 | 31.6 | 23.0 | 14.9 | 7.5 | 23.3 |
STALE [20] | ✗ | ✗ | ✓ | 40.5 | 32.3 | 23.5 | 15.3 | 7.6 | 23.8 | |
MMPrompt [9] | ✗ | ✗ | ✓ | 46.3 | 39.0 | 29.5 | 18.3 | 8.7 | 28.4 | |
ZEETAD [23] | ✗ | ✗ | ✓ | 61.4 | 53.9 | 44.7 | 34.5 | 20.5 | 43.2 | |
EffPrompt† | ✓ | ✗ | ✓ | 7.1 | 5.9 | 4.5 | 3.4 | 2.2 | 4.6 | |
STALE† | ✓ | ✗ | ✓ | 0.5 | 0.3 | 0.2 | 0.2 | 0.2 | 0.3 | |
T3AL0 [12] | ✓ | ✓ | ✓ | 11.1 | 6.5 | 3.2 | 1.5 | 0.6 | 4.6 | |
T3AL | ✓ | ✓ | ✗ | 19.2 | 12.7 | 7.4 | 4.4 | 2.2 | 9.2 | |
ZEAL-Pixtral | ✓ | ✓ | ✓ | 18.2 | 11.9 | 7.0 | 3.9 | 1.4 | 8.5 | |
ZEAL-Pixtral✛ | ✓ | ✓ | ✓ | 27.4 | 17.7 | 10.5 | 5.6 | 1.9 | 12.6 | |
ZEAL-LLaVA-OV | ✓ | ✓ | ✓ | 22.1 | 16.1 | 11.0 | 5.7 | 3.0 | 11.6 | |
ZEAL-LLaVA-OV✛ | ✓ | ✓ | ✓ | 35.1 | 23.8 | 15.2 | 7.6 | 3.5 | 17.0 | |
Zero-shot 50% Seen, 50% Unseen | EffPrompt [8] | ✗ | ✗ | ✓ | 37.2 | 29.6 | 21.6 | 14.0 | 7.2 | 21.9 |
STALE [20] | ✗ | ✗ | ✓ | 38.3 | 30.7 | 21.2 | 13.8 | 7.0 | 22.2 | |
MMPrompt [9] | ✗ | ✗ | ✓ | 42.3 | 34.7 | 25.8 | 16.2 | 7.5 | 25.3 | |
ZEETAD [23] | ✗ | ✗ | ✓ | 45.2 | 38.8 | 30.8 | 22.5 | 13.7 | 30.2 | |
EffPrompt† | ✓ | ✗ | ✓ | 5.4 | 4.4 | 3.5 | 2.7 | 1.9 | 3.6 | |
STALE† | ✓ | ✗ | ✓ | 1.3 | 0.7 | 0.6 | 0.6 | 0.4 | 0.7 | |
T3AL0 [12] | ✓ | ✓ | ✓ | 11.4 | 6.8 | 3.5 | 1.7 | 0.6 | 4.8 | |
T3AL | ✓ | ✓ | ✗ | 20.7 | 14.3 | 8.9 | 5.3 | 2.7 | 10.4 | |
ZEAL-Pixtral | ✓ | ✓ | ✓ | 17.0 | 10.9 | 6.3 | 3.5 | 1.2 | 7.8 | |
ZEAL-Pixtral✛ | ✓ | ✓ | ✓ | 26.6 | 16.7 | 9.7 | 5.1 | 1.7 | 11.9 | |
ZEAL-LLaVA-OV | ✓ | ✓ | ✓ | 21.1 | 15.0 | 9.9 | 5.0 | 2.6 | 10.7 | |
ZEAL-LLaVA-OV✛ | ✓ | ✓ | ✓ | 33.8 | 22.4 | 14.0 | 6.8 | 3.0 | 16.0 |
@@
4.2 Filtering Action Classes for Efficiency
To reduce the number of LVLM queries needed to localize actions, we filter relevant action classes to the video of interest using CLIP-ranking of action classes. Specifically, we use EVA-02-CLIP-L/14+ [28] to encode a specific number of frames uniformly sampled throughout video and measure the cosine similarity between the mean image features and embedded action class names.
We record the top- action classes and only use the and relevant to these action classes when querying LVLM. For THUMOS14, we find that is sufficient for high recall in assessing the correct action classes that could be present in video. Figure 3 depicts precision curves and recall curves for different number of frames sampled per video. From this, we also find 8 frames suffices for accurately representing video for this step of our framework.
@@ 4.3 Zero-shot Action Localization
Our method outperforms T3AL’s best baseline up to a 26.1% improvement in average mAP on the 75% — 25% split, and up to a 2.9% improvement on the 50% — 50% split, using only pretrained LVLM to affirm visual questions and generate soft-scores, and without parameter updates of any kind. If video-level labels are available (if the set of action classes for a given video is constrained), our method enjoys strong performance gains. We also note that using LLaVA-OneVision as the LVLM in our setup performs marginally better than using Pixtral as LVLM.
@@ 4.4 Ablations
In addition, we conduct ablations to investigate the contributions of different components of our framework. All ablations reported here are performed on the THUMOS14 benchmark (50% — 50% split) and assuming action classes for a given video are known.
CLIP | Average mAP |
SigLIP [35] | 15.9 |
EVA-CLIP [28] | 17.0 |
CLIP for actionness scores. First, we ablate the backbone used for deriving actionness scores for intervals (as in Section 3.3). We test EVA-02-CLIP-L/14+ [28] and SigLIP-SoViT-400M/14 [35] as strong CLIP models for image and text embedding. We show downstream localization performance in Tab. 2 when using either backbone for actionness, and we find that EVA-CLIP is the best performing CLIP backbone for our framework.
Effect of actionness scores for suppression. We also ablate the use of actionness scores via the tradeoff parameter. corresponds to where only is used to score intervals, while indicates that is ignored entirely (i.e. only boundary confidence scores contribute to an interval’s score calculation). Figure 5 depicts the effect on average mAP for varying values of the parameter, keeping all other components of the framework fixed. We observe that is the optimal value for the THUMOS14 benchmark, but also that average mAP is greater when than when . This implicates two important findings: 1) neither or alone is sufficient for assessing proposal likelihood (having highly confident boundary predictions and highly confident semantic conent within the interval is not mutually exclusive) and 2) boundary confidence is more critical to performance than CLIP-based actionness within the interval.
Limitations and Future Work. The primary limitation of the proposed framework is the computational overhead of inference with LVLM. This may impact applications where near real-time inference speeds are necessary for utility. However, in many applications, zero-shot localization is a retrospective tool for large-scale video analysis, obviating the need for fast inference. Another limitation of this framework lies in the need to filter relevant action classes to a given sample. When the number of potential action classes is large, this can limit the performance of our framework relying on action-specific queries to LVLM. One could instead use LVLM to generate captions or action classes from key frames of the video itself to limit the size of the query space, and we leave this to future work.
In addition, the reliance of our method on the logits of LVLM requires public access to the model. Also, extracting the logits of the ”yes” and ”no” tokens as described in this work may not be the only way to assess confidence of LVLM in answering visual questions. Instead, one could inflate action classes with key phrases that describe different phases of the action and observe the enrichment of the phrase’s token logits at each frame. This strategy and other prompting strategies for LLM could result in more fine-grained localization outputs, but we leave this as a promising direction for future work.
5 Conclusions